OpenStudy (anonymous):

Need help solving: 1/by the square root 2-x + square of 2-x =2, confused.

OpenStudy (anonymous):

\[\frac{1}{\sqrt{2-x}}+(2-x)^2=2\] like that?

OpenStudy (anonymous):

After the + it square just like the left side of the equation?